# Viscoelastic Worthington jets & droplets produced by bursting bubbles

### v1

@@ 3.1 Polymeric liquids with permanent memory
We stress that in this limit, the jet breakup occurs due to finite grid resolution in our numerical code (Lohse & Villermaux, 2020; Chirco et al., 2022; Kant et al., 2023). We cannot differentiate between a case of drop detachment from the jet or the case when they are still connected through a thin filament–also known as the beads-on-a-string structure (Hosokawa et al., 2023; Clasen et al., 2006; Pandey et al., 2021; Zinelis et al., 2023). Although current simulations fully resolve other aspects, they cannot resolve these finest threads, which may have subgrid cell sizes depending on the $Ec$. At higher grid resolutions, we expect to recover the beads-on-a-string configuration, as the Oldroyd-B model does not yield a finite time breakup singularity in the infinite $De$ regime, instead converging to a finite filament (Eggers et al., 2020; Turkoz et al., 2018, 2021). To prevent infinite thread thinning, a nonlinear elastic model could also be employed (see § 2.1 for further discussions).
@@
The emerging Worthington jet may eject multiple droplets. For Newtonian liquids, exact solutions for the first droplets’ size $r_{d}$ are well understood (see appendix A and Blanco-Rodríguez & Gordillo (2020)). $r_{d}$ decreases with $Oh_{s}$ until $Oh_{s}\approx 0.0375$, beyond which the droplet breaks from the jet due to the Rayleigh–Plateau instability and falls downwards. Our analysis focuses on droplets propagating away from the source, excluding those with downward velocity upon breakup (observed in Newtonian media for $0.0450.0375<Oh_{s}<$ 0.045). For elastic cases, despite unresolved filaments connecting droplets and jets, we have rigorously verified the convergence of the first droplet’s size to at least 10% accuracy. Figure 4(b) illustrates a phase map of the first droplet’s size $r_{d}$, revealing intriguing differences from the jet behavior. While $r_{d}$ follows the same trend with $Oh_{s}$ and remains invariant of $Ec$ below critical values, the critical $Ec$ for droplet suppression differs from that of jet suppression. As the jet width is determined solely by $Oh_{s}$, independently of $Ec$, the first emerging droplet’s size also remains independent initially. However, as $Ec$ increases further, rising elastic stresses suppress droplet formation more abruptly than jet formation. The critical values $Ec_{d}$ for the transition between jet formation with and without droplet breakup (dropping transition) are sensitive to $Oh_{s}$, with the critical $Ec_{d}$ decreasing as $Oh_{s}$ increases. This trend is in stark contrast with the transition from jet formation to jet suppression (jetting transition), where critical $Ec$ values remain largely $Oh_{s}$-independent.
@@ 3.2 Polymeric liquids with poor memory
To further interpret the jetting dynamics and drop formation, figure 6 presents phase maps illustrating the behavior of maximum jet lengths ($L_{\text{max}}$) and first droplet sizes ($r_{d}$) for $De=0.01$. Figure 6(a) shows $L_{\text{max}}$ across a range of $Ec$ and $Oh_{s}$ values. For low $Ec$, $L_{\text{max}}$ shows Newtonian-like $Oh_{s}$ dependence. As $Ec$ increases, $L_{\text{max}}$ decreases monotonically until jet formation ceases beyond an $Oh_{s}$-independent critical $Ec_{j}$, mirroring the infinite $De$ limit behavior. Figure 6(b) maps the $r_{d}$, showing $Ec$-independent droplet sizes until near the transition point, where droplet formation is suppressed. For $De\ll 1$, the critical $Ec_{d}$ for the dropping transition exhibits minimal $Oh_{s}$-dependence, contrasting with the $Oh_{s}$-sensitive behavior at infinite $De$. Comparing these results to the $De\to\infty$ limit reveals persistent fundamental regimes across different $De$ values, but the transition thresholds are highly sensitive to the polymeric liquid’s relaxation time. Critical $Ec$ values for both jet and droplet suppression are significantly higher at low $De$ compared to the infinite $De$ limit, indicating that rapid relaxation of polymeric stresses allows jet and droplet formation at higher $Ec$ values. This low $De$ behavior suggests an interplay between elastic and viscous effects, explored further in § 4.
@@ 4 Regime Map
The bursting bubble dynamics in viscoelastic media exhibit distinct behavior compared to Newtonian fluids. While capillary wave propagation remains largely unaffected by viscoelasticity, the Worthington jet dynamics are highly sensitive to both the elastocapillary number $Ec$ and Deborah number $De$. This sensitivity manifests in three distinct regimes: (i) jets forming droplets, (ii) jets without droplet formation, and (iii) absence of jet formation. By extending our analysis across the entire $Ec$-$De$ phase space, we elucidate the transitions between these regimes, building upon our previous examination of the limiting cases $De\to\infty$ and $De\to 0$ from § 3.
@@ 4.1 Summary of the different regimes
Figure 7 depicts the elastocapillary-Deborah number ($Ec$-$De$) phase diagram, distinguishing between jets forming droplets and those without droplet formation. Figure 8 further delineates the transition to the absence of jet formation. These transitions exhibit distinct characteristics at two asymptotic limits: $De\to\infty$ and $De\to 0$. For polymeric liquids with long relaxation times ($De\gg 1$):
@@
    the dropping transition occurs at $Ec_{d}(Oh_{s})$, exhibiting strong dependence on $Oh_{s}$ (Figures 4b and 7), and
@@
    the jetting transition takes place at $Ec_{j}\approx 0.086$, independently of $Oh_{s}$ (Figure 8a).
Notably, both transitions at $De>1$ occur asymptotically close to their corresponding solutions at $De\to\infty$, reflecting the finite process duration. When the relaxation time exceeds the process timescale, polymers lack sufficient time to relax, rendering the process $De$-independent.
For polymeric liquids with short relaxation times ($De\ll 1$), both transitions are $Oh_{s}$-independent and occur at constant polymeric Ohnesorge number $Oh_{p}=Ec\times De$:
@@
These distinct transition behaviors reveal that for very small $De$, the addition of polymers can be considered to have an added viscous effect characterized by the polymeric Ohnesorge number $Oh_{p}$. In contrast, at high $De$, the medium behaves analogously to an elastic solid for $Oh_{s}\to 0$ and as a Kelvin-Voigt solid for finite $Oh_{s}$. We further investigate the jetting transition using slender jet equations in § 4.2 following similar approaches by Driessen et al. (2013); Gordillo et al. (2020); Zinelis et al. (2023); Sen et al. (2024).
Figure 9: Temporal evolution of the Worthington jet for a representative case, where the jet emerges, reaches a maximum, and is pulled to merge with the liquid bath. The control volume contains the jet region, as shown by the region within the gray lines. Here, $h(z,t)$ is the width of the jet, which becomes $h_{\text{base}}$ at the base of the jet. The capillary force at the jet base is $F_{\gamma}=\gamma\left(2\pi h_{\text{base}}\right)$ that acts radially outwards. At the same time, the elastic and viscous stresses act at the base of the jet as ${{\color[rgb]{0,0,0}F_{\eta}+F_{p}=\left(\sigma_{\eta,\text{base}}+\sigma_{p,\text{base}}\right)\pi h_{\text{base}}^{2}}}$.
@@ 4.2 What sets the different transitions, and what do we learn from these transitions?
To understand the mechanisms governing bubble cavity collapse, we analyze jet dynamics using a control volume approach (figure 9). Employing the slender jet approximation (Eggers & Fontelos, 2015, p. 240), given the small radial-to-axial length scale ratio, the vertical momentum equation for the jet reads
@@
$\frac{d\mathcal{M}_{\text{jet}}}{dt}=3\eta_{s}h^{2}\frac{\partial v}{\partial z}\Bigg{|}_{\text{base}}+Gh^{2}(\mathcal{A}_{zz}-1)\Bigg{|}_{\text{base}}={{\color[rgb]{0,0,0}\left(\sigma_{\eta,\text{base}}+\sigma_{p,\text{base}}\right)\pi h_{\text{base}}^{2}}}$ (17)
@@ 4.2.1 The limit of $De\to\infty$
Figure 11: (a) Evolution of the maximum elastic stress at the jet base ($\text{max}\left(\sigma_{p,\text{base}}(t)\right)$), normalized by the Newtonian-like viscous stress $\sigma_{N,\text{base}}$ with viscosity $\eta_{p}=G\lambda$, as a function of $De$ for different $Oh_{p}$ at $Oh_{s}=0.001$. The gray dashed horizontal line represents $\text{max}\left(\sigma_{p,\text{base}}(t)\right)\approx\sigma_{N,\text{base}}$ while the black dashed line serves as a guide to the eye representing $\text{max}\left(\sigma_{p,\text{base}}(t)\right)/\sigma_{N,\text{base}}\sim De$. (b) The variation of jet’s tip velocity $V_{\text{jet}}$, normalized by the inertiocapillary velocity $u_{\gamma}=\sqrt{\gamma/\rho_{s}R_{0}}$, with $Oh_{\text{effective}}=3Oh_{s}+2Oh_{p}$ at different $De$ and $Oh_{s}=0.01$. The gray dashed line represents $V_{\text{jet}}\sim\gamma/\eta_{\text{effective}}$.
@@ 4.2.2 The limit of $De\to 0$
In the limit of $De\to 0$, polymeric liquids exhibit an added viscous effect characterized by the polymeric Ohnesorge number $Oh_{p}$ (also see § 4.1). To quantify this stress, we evaluate the maximum elastic stress $\text{max}(\sigma_{p,\text{base}}(t))$ normalized by a Newtonian-like viscous stress $\sigma_{N,\text{base}}$, calculated when the elastic stress reaches its maximum using
$\sigma_{N,\text{base}}=\frac{2}{h_{\text{base}}^{2}}\int_{o}^{h_{\text{base}}}G\lambda\frac{\partial v}{\partial z}h\mathrm{d}h,$ (19)
where $\eta_{p}=G\lambda$ represents the polymeric contribution to the shear viscosity of the polymeric liquid. Figure 11(a) demonstrates that $\text{max}(\sigma_{p,\text{base}}(t))$ collapse for all $Oh_{p}$ when rescaled with $\sigma_{N,\text{base}}$ for $De\to 0$.
As $De$ approaches unity, marking the onset of the infinite $De$ asymptotic regime, we observe a scaling behavior of $\text{max}(\sigma_{p,\text{base}}(t))\sim De\times\sigma_{N,\text{base}}$. This scaling remarkably resembles that predicted by Boyko et al. (2024) for flow in a slowly varying contraction at the infinite $De$ asymptote, despite significant geometric differences. While our study focuses on free surface flows and Boyko et al. (2024) examined contraction geometries, this unexpected similarity hints at a potentially universal behavior near the infinite $De$ asymptote. To further examine this intriguing connection, a similar closed-form $De$ expansion for free surface flows is necessary. However, we caution that this scaling approach to the infinite $De$ asymptote could be system-dependent (Hinch et al., 2024).
Here, we focus at the zero $De$ limit, where the elastic stress can be replaced with a Newtonian-like viscous stress with polymeric viscosity $\eta_{p}$ such that the jetting transition occurs at constant polymeric $Oh_{p}$. The polymeric stress contribution becomes $\sigma_{p}\approx 2G\lambda\boldsymbol{\mathcal{D}}$ (for the Oldroyd-B rheology, see also equation (13)). Consequently, we can rewrite the force balance in equation (17) as
$\frac{d{{\color[rgb]{0,0,0}\mathcal{M}}}_{\text{jet}}}{dt}=\left(3\eta_{s}+2G\lambda\right)h^{2}\frac{\partial v}{\partial z}\Bigg{|}_{\text{base}}{{\color[rgb]{0,0,0},}}$ (20)
which depicts the balance of jet inertia with viscous forces. Using as relevant scales for the jet momentum $\mathcal{M}_{\text{jet}}\sim\rho_{s}V_{\text{jet}}h_{\text{base}}^{3}$, where $V_{\text{jet}}$ is jet velocity, the velocity gradient $\partial_{z}v\sim V_{\text{jet}}/\delta_{\eta}$, and the time scale $\tau_{i}\sim h_{\text{base}}/V_{\text{jet}}$, we obtain the balance of forces as
$\rho V_{\text{jet}}^{2}\sim\eta_{\text{effective}}\frac{V_{\text{jet}}}{\delta_{\eta}}{{\color[rgb]{0,0,0}.}}$ (21)
Here, $\delta_{\eta}$ is the relevant length scale for viscous forces, and
$\eta_{\text{effective}}=3\eta_{s}+2G\lambda$ (22)
is the effective viscosity. As polymers do not affect the flow before jet formation (§ 3), the jet Weber number remains constant at inception (Gordillo & Rodríguez-Rodríguez, 2019),
@@
which has the same form as for Newtonian media (Gordillo & Rodríguez-Rodríguez, 2019), albeit with a modified viscosity term. This scaling holds in the viscous regime for Newtonian liquids (Blanco-Rodríguez & Gordillo, 2020).
Figure 12: The capillary waves focus and collapse at the bottom of the cavity. (a) The inception of the jet after the collapse at different $Oh_{p}$ at $De=0.001$ and $Oh_{s}=0.01$. The radius of the jet at the base $R_{\text{jet}}$ decreases with $Oh_{p}$ until $Oh_{p,c}=0.017$, beyond which bubbles are entrained. (b) Radius of jet $R_{\text{jet}}$ with $Oh_{p}$ at $De=0.001$ and $Oh_{s}=0.01$. $R_{\text{jet}}$ remains close to the value at the Newtonian limit $Oh_{p}=0$, and decreases sharply as it approaches $Oh_{p,c}$. Beyond $Oh_{p,j}$ jets are no longer observed.
@@
To further quantify this behavior, figure 12(a) illustrates jet features at inception for different $Oh_{p}$, while figure 12(b) shows jet radius as a function of $Oh_{p}$ at different $De$. At small $Oh_{p}$, the jet radius remains constant and close to the corresponding Newtonian liquid value (figure 12a: $Oh_{p}=0,0.005$). As $Oh_{p}$ increases, the jet width decreases sharply, vanishing at $Oh_{p,c}\approx 0.017$ (figure 12a: $Oh_{p}=0.013,0.018$), where $Oh_{p,c}$ marks the lower bound in $Oh_{p}$ for bubble entrainment inside the liquid (figure 12a: $Oh_{p}=0.018,0.025$ Gordillo & Rodríguez-Rodríguez (2019); Blanco-Rodríguez & Gordillo (2020); Rodríguez-Díaz et al. (2023)). Interestingly, the prediction for Newtonian liquids applies well to viscoelastic liquids by substituting $Oh_{p}$ for $Oh_{s}$ (figure 12b). Beyond $Oh_{p,c}$, the jet radius becomes ill-defined as the jet gradually widens (figure 12a: $Oh_{p}=0.035$), first reaching the dropping transition at $Oh_{p,d}\approx 0.048$ (figures 7 and 12b) and ultimately vanishing at $Oh_{p,j}\approx 0.129$ (figure 8b).

### v2

@@ 3.1 Polymeric liquids with permanent memory
We stress that in this limit, the jet breakup occurs due to finite grid resolution in our numerical code (Lohse & Villermaux, 2020; Chirco et al., 2022; Kant et al., 2023). We cannot differentiate between a case of drop detachment from the jet or the case when they are still connected through a thin filament–also known as the beads-on-a-string structure (Hosokawa et al., 2023; Clasen et al., 2006; Pandey et al., 2021; Zinelis et al., 2024). Although current simulations fully resolve other aspects, they cannot resolve these finest threads, which may have subgrid cell sizes depending on the $Ec$. At higher grid resolutions, we expect to recover the beads-on-a-string configuration, as the Oldroyd-B model does not yield a finite time breakup singularity in the infinite $De$ regime, instead converging to a finite filament (Eggers et al., 2020; Turkoz et al., 2018, 2021). To prevent infinite thread thinning, a nonlinear elastic model could also be employed (see § 2.1 for further discussions).
@@
The emerging Worthington jet may eject multiple droplets. For Newtonian liquids, predictions   for the first droplets’ size $r_{d}$ are well understood (see appendix A and Gañán-Calvo (2017); Blanco-Rodríguez & Gordillo (2020)). $r_{d}$ decreases with $Oh_{s}$ until $Oh_{s}\approx 0.0375$, beyond which the droplet breaks from the jet due to the Rayleigh–Plateau instability and falls downwards. Our analysis focuses on droplets propagating away from the source, excluding those with downward velocity upon breakup (observed in Newtonian media for $0.0450.0375<Oh_{s}<$ 0.045). For elastic cases, despite unresolved filaments connecting droplets and jets, we have rigorously verified the convergence of the first droplet’s size to at least 10% accuracy. Figure 4(b) illustrates a phase map of the first droplet’s size $r_{d}$, revealing intriguing differences from the jet behavior. While $r_{d}$ follows the same trend with $Oh_{s}$ observed at Newtonian limits   and remains invariant of $Ec$ below critical values, the critical $Ec$ for droplet suppression differs from that of jet suppression. As the jet width is determined solely by $Oh_{s}$, independently of $Ec$, the first emerging droplet’s size also remains independent initially. However, as $Ec$ increases further, rising elastic stresses suppress droplet formation more abruptly than jet formation. The critical values $Ec_{d}$ for the transition between jet formation with and without droplet breakup (dropping transition) are sensitive to $Oh_{s}$, with the critical $Ec_{d}$ decreasing as $Oh_{s}$ increases. This trend is in stark contrast with the transition from jet formation to jet suppression (jetting transition), where critical $Ec$ values remain largely $Oh_{s}$-independent.
@@ 3.2 Polymeric liquids with poor memory
To further interpret the jetting dynamics and drop formation, figure 6 presents phase maps illustrating the behavior of maximum jet lengths ($L_{\text{max}}$) and first droplet sizes ($r_{d}$) for $De=0.01$. Figure 6(a) shows $L_{\text{max}}$ across a range of $Ec$ and $Oh_{s}$ values. For low $Ec$, $L_{\text{max}}$ shows Newtonian-like $Oh_{s}$ dependence. As $Ec$ increases, $L_{\text{max}}$ decreases monotonically until jet formation ceases beyond an $Oh_{s}$-independent critical $Ec_{j}$, mirroring the infinite $De$ limit behavior. Figure 6(b) maps the $r_{d}$, showing $Ec$-independent droplet sizes that are equal to values at the Newtonian limit , until near the transition point, where droplet formation is suppressed. For $De\ll 1$, the critical $Ec_{d}$ for the dropping transition exhibits minimal $Oh_{s}$-dependence, contrasting with the $Oh_{s}$-sensitive behavior at infinite $De$. Comparing these results to the $De\to\infty$ limit reveals persistent fundamental regimes across different $De$ values, but the transition thresholds are highly sensitive to the polymeric liquid’s relaxation time. Critical $Ec$ values for both jet and droplet suppression are significantly higher at low $De$ compared to the infinite $De$ limit, indicating that rapid relaxation of polymeric stresses allows jet and droplet formation at higher $Ec$ values. This low $De$ behavior suggests an interplay between elastic and viscous effects, explored further in § 4.
@@ 4 Regime Map
The bursting bubble dynamics in viscoelastic media exhibit distinct behavior compared to Newtonian fluids. Our analysis reveals three well-defined regimes: (i) jets that form droplets, (ii) jets without droplet formation, and (iii) complete suppression of jets. While viscoelasticity significantly modifies jet dynamics, the capillary wave propagation prior to jet formation remains remarkably unaffected. This section explores the transitions between these regimes across the $Ec$-$De$ phase space, extending our earlier analysis of the limiting cases $De\to\infty$ and $De\to 0$ from § 3.
@@ 4.1 Summary of the different regimes
The transitions between these regimes depend on both $Ec$ and $De$, exhibiting markedly different characteristics in two limiting cases: $De\to\infty$ and $De\to 0$. Figure 7 maps these transitions in the elastocapillary-Deborah number ($Ec$-$De$) phase space, delineating the boundaries between droplet-forming jets and jets without droplets. Figure 8 complements this by illustrating the transition to complete jet suppression. Notably, the infinite $De$ asymptotic behavior extends down to $De\approx 1$, reflecting that polymers lack sufficient time to relax when relaxation times exceed the process timescale.
For polymeric liquids with long relaxation times ($De\gg 1$), we observe that:
@@
    the dropping transition occurs at $Ec_{d}(Oh_{s})$, with strong $Oh_{s}$ dependence (Figures 4b and 7), and
@@
    the jetting transition occurs at $Ec_{j}\approx 0.086$, independent of $Oh_{s}$ (Figure 8a).
Conversely, for polymeric liquids with short relaxation times ($De\ll 1$), we find that both transitions are $Oh_{s}$-independent and occur at constant polymeric Ohnesorge number $Oh_{p}=Ec\times De$:
@@
These behaviors reflect fundamentally different physical mechanisms: at high $De$, depending on $Oh_{s}$, the medium behaves like an elastic solid ($Oh_{s}\to 0$) or Kelvin-Voigt solid (finite $Oh_{s}$). However, at low $De$, polymer addition manifests as an enhanced viscous effect characterized by $Oh_{p}$. We further investigate the jetting transition using slender jet equations in § 4.2 following similar approaches by Driessen et al. (2013); Gordillo et al. (2020); Zinelis et al. (2024); Sen et al. (2024).
Figure 9: Temporal evolution of the Worthington jet for a representative case, where the jet emerges, reaches a maximum, and is pulled to merge with the liquid bath. The control volume contains the jet region, as shown by the region within the gray lines. Here, $h(z,t)$ is the width of the jet, which becomes $h_{\text{base}}$ at the base of the jet. The capillary force at the jet base is $F_{\gamma}=\gamma\left(2\pi h_{\text{base}}\right)$ that acts radially outwards. At the same time, the elastic and viscous stresses act at the base of the jet as ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F_{\eta}+F_{p}=\left(\sigma_{\eta,\text{base}}+\sigma_{p,\text{base}}\right)\pi h_{\text{base}}^{2}}}$.
@@ 4.2 What sets the different transitions, and what do we learn from these transitions?
To understand the mechanisms governing bubble cavity collapse, we analyze jet dynamics using a control volume approach (figure 9). Employing the slender jet approximation (Shi et al., 1994; Eggers & Fontelos, 2015; Driessen et al., 2013), given the small radial-to-axial length scale ratio, the vertical momentum equation for the jet reads
@@
$\frac{d\mathcal{M}_{\text{jet}}}{dt}=3\eta_{s}h^{2}\frac{\partial v}{\partial z}\Bigg{|}_{\text{base}}+Gh^{2}(\mathcal{A}_{zz}-1)\Bigg{|}_{\text{base}}={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(\sigma_{\eta,\text{base}}+\sigma_{p,\text{base}}\right)\pi h_{\text{base}}^{2}}}$ (18)
@@ 4.2.1 The limit of $De\to\infty$
Figure 11: (a) Evolution of the maximum elastic stress at the jet base ($\text{max}\left(\sigma_{p,\text{base}}(t)\right)$), normalized by the Newtonian-like viscous stress $\sigma_{N,\text{base}}$ with viscosity $\eta_{p}=G\lambda$, as a function of $De$ for different $Oh_{p}$ at $Oh_{s}=0.001$. The gray dashed horizontal line represents $\text{max}\left(\sigma_{p,\text{base}}(t)\right)\approx\sigma_{N,\text{base}}$ while the black dashed line serves as a guide to the eye representing $\text{max}\left(\sigma_{p,\text{base}}(t)\right)/\sigma_{N,\text{base}}\sim De$. Note that $Ec=Oh_{p}/De$. (b) The variation of jet’s tip velocity $V_{\text{jet}}$, normalized by the inertiocapillary velocity $u_{\gamma}=\sqrt{\gamma/\rho_{s}R_{0}}$, with $Oh_{\text{effective}}=3Oh_{s}+2Oh_{p}$ at different $De$ and $Oh_{s}=0.01$. The gray dashed line represents $V_{\text{jet}}\sim\gamma/\eta_{\text{effective}}$.
@@ 4.2.2 The limit of $De\to 0$
In the zero $De$ limit, polymeric liquids exhibit additional viscous effects characterized by the polymeric Ohnesorge number $Oh_{p}$ (also see § 4.1). The maximum elastic stress $\text{max}(\sigma_{p,\text{base}}(t))$, when normalized by the Newtonian-like viscous stress $\sigma_{N,\text{base}}$, collapses for all $Oh_{p}$ as $De\to 0$, where
$\sigma_{N,\text{base}}=\frac{2}{h_{\text{base}}^{2}}\int_{o}^{h_{\text{base}}}G\lambda\frac{\partial v}{\partial z}h\mathrm{d}h.$ (20)
As $De$ approaches unity, marking the onset of the infinite $De$ asymptotic regime, the elastic stress scales as $\text{max}(\sigma_{p,\text{base}}(t))\sim De\times\sigma_{N,\text{base}}$. This scaling remarkably resembles that predicted by Boyko et al. (2024) for flow in a slowly varying contraction at the infinite $De$ asymptote, despite significant geometric differences. While our study focuses on free surface flows and Boyko et al. (2024) examined contraction geometries, this unexpected similarity hints at a potentially universal behavior near the infinite $De$ asymptote. To further examine this intriguing connection, a similar closed-form $De$ expansion for free surface flows is necessary. However, we caution that this scaling approach to the infinite $De$ asymptote could be system-dependent (Hinch et al., 2024).
At zero $De$, the elastic stress reduces to a Newtonian-like viscous stress with polymeric viscosity $\eta_{p}$, yielding $\sigma_{p}\approx 2G\lambda\boldsymbol{\mathcal{D}}$ for Oldroyd-B rheology. The force balance in equation (18) becomes
$\frac{d\mathcal{M}_{\text{jet}}}{dt}=\left(3\eta_{s}+2G\lambda\right)h^{2}\frac{\partial v}{\partial z}\Bigg{|}{\text{base}},$ (21)
which depicts the balance of jet inertia with viscous forces. Using characteristic scales for jet momentum $\mathcal{M}_{\text{jet}}\sim\rho V_{\text{jet}}h_{\text{base}}^{3}$, velocity gradient $\partial_{z}v\sim V_{\text{jet}}/\delta_{\eta}$, and time $\tau_{i}\sim h_{\text{base}}/V_{\text{jet}}$, the force balance yields
$\rho V_{\text{jet}}^{2}\sim\eta_{\text{effective}}\frac{V_{\text{jet}}}{\delta_{\eta}}.$ (22)
Here, $\delta_{\eta}$ represents the viscous length scale and the effective viscosity is
$\eta_{\text{effective}}=3\eta_{s}+2G\lambda.$ (23)
Since polymers do not affect the flow before jet formation (§ 3), the jet Weber number remains constant at inception (Blanco-Rodríguez & Gordillo, 2021),
@@
analogous to Newtonian media but with modified viscosity (Gordillo & Rodríguez-Rodríguez, 2019; Blanco-Rodríguez & Gordillo, 2020).
Figure 12: The capillary waves focus and collapse at the bottom of the cavity. (a) The inception of the jet after the collapse at different $Oh_{p}$ at $De=0.001$ and $Oh_{s}=0.01$. The radius of the jet at the base $R_{\text{jet}}$ decreases with $Oh_{p}$ until $Oh_{p,c}=0.017$, beyond which bubbles are entrained and the jet radius increases. (b) Radius of jet $R_{\text{jet}}$ with $Oh_{p}$ at $Oh_{s}=0.01$ and different $De$. $R_{\text{jet}}$ remains close to the value at the Newtonian limit $Oh_{p}=0$, and decreases sharply as it approaches $Oh_{p,c}$. Beyond $Oh_{p,j}$ jets are no longer observed.
@@
To further quantify this behavior, figure 12(a) illustrates jet features at inception for different $Oh_{p}$ at $De=0.001$, while figure 12(b) shows jet radius as a function of $Oh_{p}$ at different $De$. At small $Oh_{p}$, we observe that the jet radius maintains a value comparable to the Newtonian reference case (figure 12a: $Oh_{p}=0,0.005$). This behavior is consistent with the $De\to 0$ limit, where polymeric additives primarily contribute enhanced effective viscosity. Since the jet radius determines the resulting drop size (Gañán-Calvo, 2017; Blanco-Rodríguez & Gordillo, 2020), this independence of jet radii in the low $Oh_{p}$ regime suggests minimal variation in droplet size distribution compared to Newtonian cases.   As $Oh_{p}$ increases, we observe a pronounced reduction in jet width until reaching $Oh_{p,c}\approx 0.017$ (figure 12a: $Oh_{p}=0.013,0.018$). At this critical value, the system transitions to a bubble entrainment regime (figure 12a: $Oh_{p}=0.018,0.025$ Gordillo & Rodríguez-Rodríguez (2019); Blanco-Rodríguez & Gordillo (2020); Rodríguez-Díaz et al. (2023)). Interestingly, the prediction for Newtonian liquids applies well to viscoelastic liquids by substituting $Oh_{p}$ for $Oh_{s}$ (figure 12b), particularly in the $De\to 0$ limit. Beyond $Oh_{p,c}$, the jet radius becomes ill-defined as the jet gradually widens (figure 12a: $Oh_{p}=0.035$), first reaching the dropping transition at $Oh_{p,d}\approx 0.048$ (figures 7 and 12b) and ultimately vanishing at $Oh_{p,j}\approx 0.129$ (figure 8b).